# On Schultz’s generalization of Borweins’ cubic identity

### v1

@@ Abstract.
Around 1991, J.M. and P.B. Borwein established a cubic analogue of Jacobi’s fundamental identity for theta functions. Their identity serves as the foundation for the subsequent development by B.C. Berndt, S. Bhargava, and F.G. Garvan of Ramanujan’s theory of elliptic functions to the cubic base. In 2013, D. Schultz discovered an identity for theta series in three variables which generalizes the Borweins’ identity. In this article, we revisit Schultz’s identity and present two distinct approaches to its derivation. Our investigation not only provides new proofs but also yields several identities of a similar type. Furthermore, this study enables us to construct new two-variable generalizations of Jacobi’s original classical identity.
2020 Mathematics Subject Classification:
33E05, 11F27
The second author is partially supported by the Ministry of Education, Singapore, Academic Research Fund, Tier 1 (RG15/23).
The third author is partially supported by the National Natural Science Foundation of China (Grant No. 12371328) and the Ministry of Education, Singapore, Academic Research Fund, Tier 1 (RG15/23).
@@ 1. Introduction
where $\omega=e^{2\pi i/3}$. Their identity is a cubic analogue of Jacobi’s classical identity [14, p.90, (14)], [13, Theorem 3.7, Corollary 3.8] given by
@@
$\kappa(q)=\left(\dfrac{\displaystyle\sum_{m,n\in\mathbf{Z}}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}}{\displaystyle\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}}\right)^{3},$
$_{2}F_{1}\left(a,b;c;z\right)=\sum_{j=0}^{\infty}\dfrac{(a)_{j}(b)_{j}}{(c)_{j}}\dfrac{z^{j}}{j!}$
@@
Borweins’ inversion formula (1.3) is crucial in proving Ramanujan’s identity [17, (31)], [5]
@@
In 2013, D. Schultz [19] published the identity given by
@@
$+\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}x^{m}y^{n}\right)^{3}=\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}}x^{m}y^{n}\right)^{3}$
@@
Schultz’s identity turns out to be a two-variable generalization of the (1.1). To see this, we need to show that
@@
Identity (1.5) is equivalent to
$S=\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}+m+n}=0.$
Replacing $(m,n)$ by (n,mn1)(n,-m-n-1), we deduce that
$S=\omega^{-1}.S$
This implies that $S=0$ and the proof of (1.5) is complete. For other proofs of (1.5), see [23, Corollary] and [20, Section 2.3.3]. Substituting $x=y=1$ in (1.4) and using (1.5), we arrive at (1.1).
The first proof of (1.4) was provided by Schultz [19, p. 622]. He also mentioned, without giving explicit details, four approaches [19, p. 663] to establish the identity. The only other proof of (1.4) which we are aware of is due to X.R. Ma and R.Z. Wei [16, Corollary 4.12].
In this article, we give two new proofs of (1.4). Our first approach is motivated by the work of X.M. Yang [24]. In Section 2, we show that (1.4) follows from one of her identities [24, (3.8)]. In Section 4, we use transformation formulas for theta functions and the identities in Sections 2 and 3 to derive analogues of (1.4).
Our second proof of (1.4) connects the identity to the study of Macdonald’s identities [9] and this is presented in Section 4.
Motivated by (1.4), we found an analogue of (1.4) for Jacobi’s identity (1.2) given by
@@
By (1.7), we see that (1.2) follows from (1.6) with the substitution $x=y=1$.
In Section 5, we derive several extensions of (1.6). In Section 6, we derive some identities equivalent to (1.1). The identities presented in these last two sections are motivated by the recent discoveries of H.H. Chan, S.H. Chan and P. Sole [8].
2. Jacobi’s theta functions and (1.4)
@@
The infinite product representations of $\vartheta_{1}(z|\tau)$ and $\vartheta_{3}(z|\tau)$ ([22, p. 469], [7, p. 30]), which are equivalent to Jacobi’s triple product identity, are given by
@@
$\Lambda=\left\{x\pi+y\pi\tau\big|0\leq x<1,0\leq y<1\right\}$
@@
which implies that $c_{2}=0$ since $\vartheta_{1}(\pi\tau|3\tau)\neq 0$. Therefore, $c_{2}=c_{3}=0$. Identity (2.6) reduces to
@@
Let
@@
The function $T(z,x,y|\tau)$ is constructed so that it satisfies (2.5) since
@@
$\vartheta_{1}(3z|3\tau)/T(z,x,y|\tau),$
$e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)/T(z,x,y|\tau)$
and
$e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau)/T(z,x,y|\tau)$
are all elliptic functions with periods $\pi$ and $\pi\tau$ and have 3 poles in $\Lambda$. Since an elliptic function with a simple pole must be a constant, we conclude that there exists $c,\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ which are dependent on $x$ and $y$ but independent of $z$ such that
@@
$=\lambda_{1}\vartheta_{1}(3z|3\tau)+\lambda_{2}e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+\lambda_{3}e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau).$ (2.8)
@@
$\qquad\qquad=-\lambda_{1}\vartheta_{3}(3z|3\tau)+\lambda_{2}e^{2iz}\vartheta_{3}(3z+\pi\tau|3\tau)+\lambda_{3}e^{-2iz}\vartheta_{3}(3z-\pi\tau|3\tau).$ (2.9)
@@
The identities we derived are recorded as follow:
Theorem 2.1.
Let $z,x,y\in\mathbf{C}$ and $q=e^{\pi i\tau}$ with $\text{Im}\,\tau>0$. Let
and
Then
$\quad+q^{1/3}C(-x,-y|\tau)e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+q^{1/3}C(x,y|\tau)e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau),$
and
We are now ready to prove (1.4).
Proof of (1.4).
@@
The proof of (2.16) using (2.15) is inspired by the proof of Theorem 2.3 in [6]. The final step of proving (1.4) using the fact that the product on the right hand side of (2.16) can be written as a difference of the cube of two theta series in two ways is inspired by the proof of Corollary 4.2 in [16].
3. Analogues of (1.4)
In this section, we derive several identities analogous to (1.4). By setting $z=0$ in (2.12) and simplifying using (2.14), we deduce that
@@
$C(x,y|\tau)-\omega^{2}C(-x,-y|\tau)=\frac{-i\omega}{\eta(\tau)}\vartheta_{1}\left(\frac{\pi}{3}+x|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}+y|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}-x-y|\tau\right).$ (3.2)
@@
The right-hand side of (3.6) is invariant under the change $x\to x+\pi/3$ and therefore, the following new analogue of (2.17) holds:
Theorem 3.1.
Let $C(x,y|\tau)$ be given by (2.11). Then
@@
Using the definition $C(x,y|\tau)$ given by (2.11) and simplifying, we deduce that (3.8) is equivalent to
$\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2})}\right)^{3}-\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n+1/3)}\right)^{3}=\frac{\eta^{9}(\tau)}{\eta^{3}(3\tau)}.$ (3.9)
@@
Equating the constant terms on both sides, we find that
@@
For an elegant proof of (3.11), see [6, Proposition 2.2(i)].
@@
Let $x=\pi/3$ in (3.12) and using
$\vartheta_{1}(\pi/3|\tau)=\sqrt{3}\eta(3\tau),$
we deduce that
$\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n+1/3)}=3\frac{\eta^{3}(3\tau)}{\eta(\tau)}.$
Setting $x=y=\pi/9$ in (3.6) and again using
$\vartheta_{1}(\pi/3|\tau)=\sqrt{3}\eta(3\tau),$
we deduce that
$\omega^{2}\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n)}\omega^{m+n}\right)^{3}-\omega^{-2}\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n)}\omega^{-m-n}\right)^{3}$
$=-3\sqrt{3}i\prod_{n=1}^{\infty}\frac{(1-q^{2n})^{6}(1-q^{18n})^{3}}{(1-q^{6n})^{3}}.$ (3.13)
@@
In the remaining of this section, we continue our search for identities similar to (1.4). Recall from [18, (5), p. 205] that if $M$ is a real symmetric $f\times f$ matrix, $\mathbf{m}\in\mathbf{Z}^{f}$ written as a column vector, $\mathbf{x}$ is a column vector with $f$ complex variable components $x_{1},x_{2},\cdots,x_{f}$, then for $\text{Im}\,\tau>0$, we have
@@
Using (3.17), (3.20) and the transformation formula [22, p. 475], [21, p. 35]
for $j=1,3$, we deduce the following identities from Theorem 2.1.
@@
4. Second proof of (1.4)
We will now give a second proof of (2.15), which leads to a second proof of (1.4).
Second proof of (2.15).
@@
$\{(m,n)|m,n\in\mathbf{Z}\}=\{(k+\ell,k-\ell)|k,\ell\in\mathbf{Z}\}\cup\{(k+\ell+1,k-\ell)|k,\ell\in\mathbf{Z}\}.$
Let $\gamma=1/2+\sqrt{-3}/2$. Note that $\gamma$ is a sixth-root of unity and $\gamma^{2}=\omega$. Let
@@
$=-e^{4iu}q^{2/3}\dfrac{i}{\eta(\tau)}\vartheta_{1}(2u-2\pi\tau/3|\tau)\vartheta_{1}(u+v-\pi\tau/3|\tau)\vartheta_{1}(u-v-\pi\tau/3|\tau).$ (4.3)
We will now show that (4.3) follows from an identity found by H.H. Chan, S. Cooper and P.C. Toh [9, Theorem 3.1] given by
@@
$G_{2}(2z|2\tau)=2\sum_{\alpha\equiv 1\!\!\!\!\!\!\pmod{6}}\!\!\!\!\!\!q^{\alpha^{2}/6}\sin{2\alpha z}$
@@
$G_{3}(2z|2\tau)=2\sum_{\alpha\equiv-2\!\!\!\!\!\!\pmod{6}}\!\!\!\!\!\!q^{\alpha^{2}/6}\sin{2\alpha z}.$
Using $\sin{u}=\dfrac{e^{iu}-e^{-iu}}{2i}$, we find that
@@
5. Proof of (1.6) and its analogues
Recently, Chan, Chan and Sole discover the identities [8, (1.5),(1.6)],

### v2

@@ Abstract.
In 1991, the Borweins established a cubic analogue of Jacobi’s identity for theta functions, which is used by B.C. Berndt, S. Bhargava, and F.G. Garvan in the development of Ramanujan’s cubic theory of elliptic functions. In 2013, D. Schultz discovered an identity for theta series in three variables which generalizes the Borweins’ identity. In this article, we revisit Schultz’s identity and present two distinct approaches to its derivation. Our investigation not only provides new proofs but also yields several new Schultz-type identities.
@@ 1. Introduction
where $\omega=e^{2\pi i/3}$. Their identity is a cubic analogue of Jacobi’s classical identity [17, p. 90, (14)], [16, Theorem 3.7, Corollary 3.8] given by
@@
Let
$_{2}F_{1}\left(a,b;c;z\right)=\sum_{j=0}^{\infty}\dfrac{(a)_{j}(b)_{j}}{(c)_{j}}\dfrac{z^{j}}{j!},$
where
@@
$\kappa(q)=\dfrac{c^{3}(q)}{a^{3}(q)}$
with
$c(q)=\sum_{m,n\in\mathbf{Z}}q^{(m+1/3)^{2}+(m+1/3)(n+1/3)+(n+1/3)^{2}}$
@@
$a(q)=\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}.$
Borweins’ inversion formula (1.3) is crucial, for example, in proving Ramanujan’s identity [21, (31)], [5]
@@
In 2013, D. Schultz [23] published a two-variable generalization of (1.1).
Theorem 1.1.
The following identity is valid:
@@
$\quad+\left(\sum_{m,n\in\mathbf{Z}}q^{m^{2}+mn+n^{2}}x^{m}y^{n}\right)^{3}=\left(\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}}x^{m}y^{n}\right)^{3}$
@@
We shall refer to (1.4) as Schultz’s identity. To see that (1.1) is a specialization of (1.4), we need to set $x=y=1$ in (1.4) and show that
@@
Identity (1.5) is equivalent to $S=0$ with
$S=\sum_{m,n\in\mathbf{Z}}\omega^{m-n}q^{m^{2}+mn+n^{2}+m+n}.$
Replacing $(m,n)$ by (n,mn1)(n,-m-n-1) in the latter double sum, we deduce that
$S=\omega^{-1}\cdot S,$
which implies that $S=0$ and completes the proof of (1.5). For other proofs of (1.5), see [27, Corollary] and [24, Section 2.3.3].
The first proof of Theorem 1.1 was provided by Schultz [23, p. 622]. On page 663 of [23], Schultz mentioned four possible methods to establishing Theorem 1.1. His proof was carried out using his first method. Schultz’s remaining three methods indicated that Theorem 1.1 may be proved by
  1. (1)
    multiplying out and combining the series on both sides of (1.4),
  2. (2)
    verifying that both sides of (1.4) agree when written in terms of some basis of the theta functions of order 3,
  3. (3)
    verifying that both sides of (1.4) agree up to a certain finite order at a finite set of points.
Schultz mentioned the difficulties of proving (1.4) using the above methods but gave little details. Around 2017, X.R. Ma and R.Z. Wei [20, Corollary 4.12] gave a new proof of Theorem 1.1 which did not follow any of Schultz’s suggestions.
In this article, we give two new proofs of Theorem 1.1 which are also independent of Schultz’s insights. Our first approach is motivated by the work of X.M. Yang [28]. In Section 2, we show that (1.4) follows from one of her identities [28, (3.8)]. In Section 4, we use transformation formulas for theta functions and the identities in Sections 2 and 3 to derive analogues of (1.4). Our second proof of Theorem 1.1 connects (1.4) to the study of Macdonald’s identities [11], and this is presented in Section 4.
Motivated by Theorem 1.1, we found an analogue of (1.4) for Jacobi’s identity (1.2).
Theorem 1.2.
We have the following identity:
@@
Identity (1.6) is not new and, presumably, was known to Jacobi; its traces can be found in classical textbooks, e.g., as part of Exercise 1 in [26, Chap. 21] or [19, p. 9].
@@
[…]
In Section 5, we present and prove several new extensions of Theorem 1.2. In Section 6, we derive further identities equivalent to (1.1). The identities presented in these last two sections are motivated by the recent discoveries of H.H. Chan, S.H. Chan and P. Sole [10].
2. Jacobi’s theta functions and the first proof of Schultz’s identity
@@
The infinite product representations of $\vartheta_{1}(z|\tau)$ and $\vartheta_{3}(z|\tau)$ (see [26, p. 469], [9, p. 30]), which are equivalent to Jacobi’s triple product identity, are given by
@@
$\Lambda=\left\{x\pi+y\pi\tau\bigm|0\leq x<1,\,0\leq y<1\right\}$
@@
which implies that $c_{2}=0$ since $\vartheta_{1}(\pi\tau|3\tau)\neq 0$. Thus, $c_{2}=c_{3}=0$. Identity (2.6) reduces to
@@
Theorem 2.1.
Let $z,x,y\in\mathbf{C}$ and $q=e^{\pi i\tau}$ with $\operatorname{Im}\tau>0$. Take
and
Then
$\quad+q^{1/3}C(-x,-y|\tau)e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+q^{1/3}C(x,y|\tau)e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau)$
and
Proof.
Introduce
@@
This function is constructed so that it satisfies (2.5), since
@@
$\frac{\vartheta_{1}(3z|3\tau)}{T(z,x,y|\tau)},\quad\frac{e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)}{T(z,x,y|\tau)}\quad\text{and}\quad\frac{e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau)}{T(z,x,y|\tau)}$
are all elliptic functions with periods $\pi$ and $\pi\tau$ and with 3 poles in $\Lambda$. Since an elliptic function with a simple pole must be a constant, we conclude that there exists $c,\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ which are dependent on $x$ and $y$ but independent of $z$ such that
@@
$\qquad=\lambda_{1}\vartheta_{1}(3z|3\tau)+\lambda_{2}e^{2iz}\vartheta_{1}(3z+\pi\tau|3\tau)+\lambda_{3}e^{-2iz}\vartheta_{1}(3z-\pi\tau|3\tau).$ (2.12)
@@
$\qquad=-\lambda_{1}\vartheta_{3}(3z|3\tau)+\lambda_{2}e^{2iz}\vartheta_{3}(3z+\pi\tau|3\tau)+\lambda_{3}e^{-2iz}\vartheta_{3}(3z-\pi\tau|3\tau).$ (2.13)
@@
Thus, the required identities (2.10) and (2.11) follow from (2.12) and (2.13). ∎
We are now ready to prove Theorem 1.1.
First proof of Theorem 1.1.
@@
The proof of (2.16) using (2.15) is inspired by the proof of Theorem 2.3 in [7]. The final step in the above proof of Theorem 1.1, using the fact that the product on the right hand side of (2.16) can be written as a difference of the cube of two theta series in two ways, is inspired by the proof of Corollary 4.2 in [20].
3. Analogues of Schultz’s identity
In this section, we derive several identities analogous to (1.4).
Theorem 3.1.
Let $C(x,y|\tau)$ be given by (2.9). Then
Proof.
By setting $z=0$ in (2.10) and simplifying using (2.14), we deduce that
@@ 3. Analogues of Schultz’s identity
$C(x,y|\tau)-\omega^{2}C(-x,-y|\tau)=\frac{-i\omega}{\eta(\tau)}\vartheta_{1}\left(\frac{\pi}{3}+x|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}+y|\tau\right)\vartheta_{1}\left(\frac{\pi}{3}-x-y|\tau\right)$ (3.3)
@@
The right-hand side of (3.7) is invariant under the change $x\to x+\pi/3$, and this leads to the required identity (3.1), which is a variant of (2.17). ∎
Corollary 3.1.
We have the identity
$\frac{\eta^{9}(\tau)}{\eta^{3}(3\tau)}=\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2})}\right)^{3}-\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n+1/3)}\right)^{3}.$ (3.8)
Proof.
@@
Using the definition $C(x,y|\tau)$ given by (2.9) and simplifying, we deduce that (3.9) is equivalent to (3.8). ∎
Corollary 3.2.
The following identity is valid:
Proof.
@@
Equating the constant terms on both sides, we obtain identity (3.10). ∎
@@
For an elegant proof of Theorem 3.2, see [7, Proposition 2.2(i)].
Corollary 3.3.
The following is valid:
$\frac{\eta^{3}(3\tau)}{\eta(\tau)}=\frac{1}{3}\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n+1/3)}.$ (3.12)
Proof.
@@
Letting $x=\pi/3$ in (3.13) and using $\vartheta_{1}(\pi/3|\tau)=\sqrt{3}\eta(3\tau)$, we arrive at (3.12). ∎
Corollary 3.4.
We have the identity
$-3\sqrt{3}i\prod_{n=1}^{\infty}\frac{(1-q^{2n})^{6}(1-q^{18n})^{3}}{(1-q^{6n})^{3}}$ $=\omega^{2}\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n)}\omega^{m+n}\right)^{3}$
$\kern-17.07164pt-\omega^{-2}\left(\sum_{m,n=-\infty}^{\infty}q^{2(m^{2}+mn+n^{2}+m+n)}\omega^{-m-n}\right)^{3}.$ (3.14)
Proof.
This follows from setting $x=y=\pi/9$ in (3.7) and once again using $\vartheta_{1}(\pi/3|\tau)=\sqrt{3}\eta(3\tau)$. ∎
@@
In the remaining of this section, we continue our search for identities similar to (1.4). Recall from [22, (5), p. 205] that if $M$ is a real symmetric $f\times f$ matrix, $\mathbf{m}\in\mathbf{Z}^{f}$ written as a column vector, $\mathbf{x}$ is a column vector with $f$ complex variable components $x_{1},x_{2},\cdots,x_{f}$, then for $\operatorname{Im}\tau>0$, we have
@@
Proof.
Using (3.18), (3.21) and the transformation formula [26, p. 475], [25, p. 35]
for $j=1,3$, we deduce the identities claimed from Theorem 2.1. ∎
4. Second proof of Schultz’s identity
We will now give a second proof of (2.15) which leads to a second proof of Theorem 1.1.
Second proof of Theorem 1.1.
@@ 4. Second proof of Schultz’s identity
$\{(m,n)\mid m,n\in\mathbf{Z}\}=\{(k+\ell,k-\ell)\mid k,\ell\in\mathbf{Z}\}\cup\{(k+\ell+1,k-\ell)\mid k,\ell\in\mathbf{Z}\}.$
Let $\gamma=\exp(\pi i/3)=(1+\sqrt{-3})/2$. Note that $\gamma$ is a sixth-root of unity and $\gamma^{2}=\omega$. Let
@@
$\quad=-e^{4iu}q^{2/3}\dfrac{i}{\eta(\tau)}\vartheta_{1}(2u-2\pi\tau/3|\tau)\vartheta_{1}(u+v-\pi\tau/3|\tau)\vartheta_{1}(u-v-\pi\tau/3|\tau).$ (4.3)
We will now show that (4.3) follows from an identity found by Chan, Cooper and Toh [11, Theorem 3.1] given by
@@
$G_{2}(2z|2\tau)=2\sum_{\alpha\equiv 1\;\allowbreak(\operatorname{mod}6)}\!\!\!\!\!\!q^{\alpha^{2}/6}\sin{2\alpha z}$
@@
$G_{3}(2z|2\tau)=2\sum_{\alpha\equiv-2\;\allowbreak(\operatorname{mod}6)}\!\!\!\!\!\!q^{\alpha^{2}/6}\sin{2\alpha z}.$
Using $\sin{u}=(e^{iu}-e^{-iu})/(2i)$, we find that
@@
5. Generalizations of Schult’z identity
Recently, Chan, Chan and Sole discover the identities [10, (1.5), (1.6)],